an investor invested a total of $1100 in two mutual funds. one fund earned a 8% profit while the other earned a 6% profit. if the investors total profit was $70, how much was invested in each mutual fund?
Answer by stanbon(75887)

an investor invested a total of $1100 in two mutual funds. one fund earned a 8% profit while the other earned a 6% profit. if the investors total profit was $70, how much was invested in each mutual fund?
--------
Interest Eq:: 0.08x + 0.06(1100-x) = 70
-------
8x + 6*1100 - 6x = 7000
---
2x = 400
x = $200 (amt. invested at 8%)
1100-x = $900 (amt. invested at 6%)
===============
Cheers,
Stan H.
---------------